# Bayesian learning with Gaussian processes for low-dimensional representations of time-dependent nonlinear systems

### v1

@@ 5 Numerical examples of GP-BayesOpInf
in which the signal variance $\sigma^{2}>0$ and the lengthscale $\ell>0$ are the kernel hyperparameters.
@@ 5.1.2 ROM structure
which is eq. 2.5 with $\hat{\mathbf{O}}=[~{}\mathbf{\hat{c}}~{}~{}\mathbf{\hat{A}}~{}~{}\mathbf{\hat{H}}~{}]\in\mathbb{R}^{r\times d(r)}$, $\mathbf{d}(\hat{\mathbf{q}})=[~{}1~{}~{}\hat{\mathbf{q}}^{\mathsf{T}}~{}~{}(\hat{\mathbf{q}}\otimes\hat{\mathbf{q}})^{\mathsf{T}}~{}]^{\mathsf{T}}\in\mathbb{R}^{d(r)}$, and $d(r)=1+r+r^{2}$. The nonquadratic terms $\hat{\mathbf{c}}\in\mathbb{R}^{r}$ and $\hat{\mathbf{A}}\in\mathbb{R}^{r\times r}$ appear due to the centering term $\bar{\mathbf{q}}$ [63]. Because the product $\hat{\mathbf{q}}\otimes\hat{\mathbf{q}}=\text{vec}(\hat{\mathbf{q}}\hat{\mathbf{q}}^{\mathsf{T}})\in\mathbb{R}^{r^{2}}$ has only $r(r+1)/2<r^{2}$ degrees of freedom due to symmetry, in practice the quadratic operator $\hat{\mathbf{H}}\in\mathbb{R}^{r\times r^{2}}$ can be fully represented by an $r\times r(r+1)/2$ matrix that acts on a compressed version of the Kronecker product (see, e.g., Appendix B of [38]).
@@ 5.1.3 Noisy training data
Figure 3 shows the POD singular value decay of the snapshot data, which decreases sharply and levels out after $6$ modes. Compressed snapshot data for mode $7$ and greater concentrate near zero and do not result in a tight GP fit, hence we set the ROM dimension to $r=6$. The GP estimation times ${\bf\it t}^{\prime}$ are defined to be $m^{\prime}=400$ uniformly spaced times in the interval $[t_{0},t_{m-1}]$. After applying Algorithm 3 to obtain a posterior distribution for the reduced operators $\hat{\mathbf{O}}$, we sample from the posterior $500$ times and solve the corresponding ROMs eq. 5.14 using the same Runge–Kutta scheme as the before. The uncertainty stemming from the data noise propagates to the ROM solutions and forms a prediction distribution. To characterize the distribution, we compute the mean and the $2.5\%$ and $97.5\%$ quantiles of the samples at each time instance, see Figure 4. The individual draws are nearly identical in the training regime, but at the start of the prediction regime the width of the distribution increases noticeably and thereafter generally continues to increase in time at a relatively slow rate.
@@ 5.1.4 Sparse training data
Next, consider a scenario in which the observed data are sparse in time. Let $m=50$, again with final observation time $t_{m-1}=0.06~{}\text{s}$ but now with only $\xi=1\%$ relative noise. This is the scenario depicted in Figure 2. Results with $500$ posterior samples are shown in Figure 9 in the reduced space and Figure 10 in the physical state space. The results are similar in quality to the previous experiments: the truth is reconstructed accurately in the training regime and the sample mean of the ROM predictions is a reasonable estimate of the truth in the prediction regime. To further probe the performance of our method on sparse data, we repeat the experiment with only $m=20$ observations and again with $m=80$ observations, all drawn uniformly over the same training time domain. Figure 11 displays the results at a single spatial location. With only $m=20$ data points, GP-BayesOpInf fails to produce accurate solutions and also does not produce a wide prediction distribution. This is partially due to not having sufficient data to perform a good GP fit, but also because Algorithm 2 prioritizes stable but inaccurate ROM solutions over ROM solutions that are accurate for a limited time but which become unstable before the final simulation time. On the other hand, using $m=80$ training points produces a slightly more accurate sample mean and tighter prediction distribution than with $m=50$, i.e., more training data leads to improved accuracy.
@@ 5.2.4 Prediction at new input parameters
Figure 13: Bayesian ROM predictions for the nonlinear diffusion-reaction problem eq. 5.15aeq. 5.15c for three values of the input parameters $(a,b)$. For each set of input parameters, data are observed at $m=20$ points over the time domain $[0,1]$ with $\xi=5\%$ relative noise. Five hundred posterior samples are used to generate ROM solutions, and the resulting distribution is characterized by the pointwise sample mean and $95\$$ interquantile range.
@@ 7 Concluding remarks
There are several avenues for future work related to GP-BayesOpInf. Our experiments used squared-exponential GP kernels and linear dimensionality reduction by POD, but the methodology is flexible and can be applied with any sufficiently differential GP kernel and with any type of compression (including nonlinear manifold approximators such as in [3, 20, 21, 32, 54]). It would be interesting to explore if embedding desirable properties in the kernel can improve model performance for problems with specific characteristics (for example, periodicity). In the presence of large amounts of noise, the number of reliable reduced modes that can be obtained through POD is limited. To reduce the projection error by including more reduced modes in GP-BayesOpInf, it may be advantageous to leverage filtering strategies as shown in [19] prior to applying dimensionality reduction. We also plan to explore more advanced GP techniques, such as sparse GPs [60, 61] to reduce the trajectory data needed for regression (i.e., selecting $m^{\prime}$ automatically), and multifidelity GPs [2, 11, 37, 42, 46, 47] to fuse data from different sources. Another promising direction is to leverage the quantified uncertainties to guide the active selection of initial conditions and/or input functions for parsimoniously generating informative training trajectories.

### v2

@@ 5 Numerical examples of GP-BayesOpInf
in which the signal variance $\sigma^{2}>0$ and the lengthscale $\ell>0$ are the kernel hyperparameters. Codes to reproduce the experiments are available at https://github.com/sandialabs/GP-BayesOpInf.
@@ 5.1.2 ROM structure
Figure 4: True reduced state time derivatives, as well as GP and finite difference estimates of the time derivatives, using $m=200$ observations with $\xi=3\%$ relative noise. The time derivative GPs have small variances and produce highly accurate estimates (left), unlike the finite difference estimates, which in some instances differ from the truth by an order of magnitude (right).
which is eq. 2.5 with $\hat{\mathbf{O}}=[~{}\mathbf{\hat{c}}~{}~{}\mathbf{\hat{A}}~{}~{}\mathbf{\hat{H}}~{}]\in\mathbb{R}^{r\times d(r)}$, $\mathbf{d}(\hat{\mathbf{q}})=[~{}1~{}~{}\hat{\mathbf{q}}^{\mathsf{T}}~{}~{}(\hat{\mathbf{q}}\otimes\hat{\mathbf{q}})^{\mathsf{T}}~{}]^{\mathsf{T}}\in\mathbb{R}^{d(r)}$, and $d(r)=1+r+r^{2}$. The nonquadratic terms $\hat{\mathbf{c}}\in\mathbb{R}^{r}$ and $\hat{\mathbf{A}}\in\mathbb{R}^{r\times r}$ appear due to the centering term $\bar{\mathbf{q}}$ [67]. Because the Kronecker product $\hat{\mathbf{q}}\otimes\hat{\mathbf{q}}=\text{vec}(\hat{\mathbf{q}}\hat{\mathbf{q}}^{\mathsf{T}})\in\mathbb{R}^{r^{2}}$ has only $r(r+1)/2<r^{2}$ degrees of freedom due to symmetry, in practice the quadratic operator $\hat{\mathbf{H}}\in\mathbb{R}^{r\times r^{2}}$ can be fully represented by an $r\times r(r+1)/2$ matrix that acts on a compressed version of the Kronecker product (see, e.g., Appendix B of [41]).
@@ 5.1.3 Noisy training data
Figure 3 shows the POD singular value decay of the snapshot data, which decreases sharply and levels out after $6$ modes. Compressed snapshot data for mode $7$ and greater concentrate near zero and do not result in a tight GP fit, hence we set the ROM dimension to $r=6$. The GP estimation times ${\bf\it t}^{\prime}$ are defined to be $m^{\prime}=400$ uniformly spaced times in the interval $[t_{0},t_{m-1}]$. To further validate the quality of the training data produced by each GP regression, we solve eq. 5.2 on a fine time grid and explicitly evaluate the right-hand side of eq. 5.2g
@@
(without adding noise) to obtain time derivatives $(\partial_{t}\rho,\partial_{t}[\rho v],\partial_{t}[\rho e])$, which are then transformed to the specific-volume formulation $(\partial_{t}v,\partial_{t}p,\partial_{t}\xi)$ and compressed using the POD basis. Figure 4 plots these true time derivatives together with the GP estimates for the time derivatives which, over the time grid ${\bf\it t}^{\prime}$ and for the $i$-th reduced state, are given by the mean vector $\tilde{\mathbf{z}}_{i}=\mathbf{K}_{i}^{zy}(\mathbf{K}_{i}^{yy})^{-1}\mathbf{y}_{i}$ with covariance $(\mathbf{W}_{i}^{zz})^{-1}=\mathbf{K}_{i}^{zz}-\mathbf{K}_{i}^{zy}(\mathbf{K}_{i}^{yy})^{-1}\mathbf{K}_{i}^{yz}$. The GP estimates are highly faithful to the true time derivatives, and the associated variances so small that the region within $10$ standard deviations of the mean is only slightly visible. As a comparison, we also apply second-order finite differences to the noisy compressed snapshots to produce time derivative estimates, the standard practice for OpInf problems. Although the POD basis filters much of the noise from the high-dimensional observations (see Figure 3), the remaining noise results in significant error in the finite difference estimates, rendering them unsuitable for populating the OpInf regression eq. 2.7.
After applying Algorithm 3 to obtain a posterior distribution for the reduced operators $\hat{\mathbf{O}}$, we sample from the posterior $500$ times and solve the corresponding ROMs eq. 5.14 using the same Runge–Kutta scheme as before. The uncertainty stemming from the data noise propagates to the ROM solutions and forms a prediction distribution. To characterize the distribution, we compute the mean and the $2.5\%$ and $97.5\%$ quantiles of the samples at each time instance, see Figure 5. The individual draws are nearly identical in the training regime, but at the start of the prediction regime the width of the distribution increases noticeably and thereafter generally continues to increase in time at a relatively slow rate.
@@
Figure 10: True reduced states, observations, and trained GPs for reduced modes 1, 3, and 5 (left), and corresponding reduced state time derivatives, GP estimates, and finite difference estimates (right).
@@ 5.1.4 Sparse training data
Next, consider a scenario in which the observed data are sparse in time. Let $m=50$, again with final observation time $t_{m-1}=0.06~{}\text{s}$ but now with only $\xi=1\%$ relative noise. This is the scenario depicted in Figure 2. Figure 10 shows GP regressions of the compressed state data and predictions for the associated time derivatives. In this case, finite differences of the noisy snapshots result in estimates for the time derivatives that are much less unwieldy than in the previous experiment (see Figure 4), but they are still significantly less accurate than the GP time derivative estimates. Additionally, the GPs yield accurate training data for both the reduced state and the time derivative over the entire training time regime, not just at the sparse observation times.
@@
Results with $500$ posterior samples are shown in Figure 11 in the reduced space and Figure 12 in the physical state space. The results are similar in quality to the previous experiments: the truth is reconstructed accurately in the training regime and the sample mean of the ROM predictions is a reasonable estimate of the truth in the prediction regime. To further probe the performance of our method on sparse data, we repeat the experiment with only $m=20$ observations and again with $m=80$ observations, all drawn uniformly over the same training time domain. Figure 13 displays the results at a single spatial location. With only $m=20$ data points, GP-BayesOpInf fails to produce accurate solutions and also does not produce a wide prediction distribution. This is partially due to not having sufficient data to perform a good GP fit, but also because Algorithm 2 prioritizes stable but inaccurate ROM solutions over ROM solutions that are accurate for a limited time but which become unstable before the final simulation time. On the other hand, using $m=80$ training points produces a slightly more accurate sample mean and tighter prediction distribution than with $m=50$, i.e., more training data leads to improved accuracy.
@@ 5.2.4 Prediction at new input parameters
Figure 15: Bayesian ROM predictions for the nonlinear diffusion-reaction problem eq. 5.15aeq. 5.15c for three values of the input parameters $(a,b)$. For each set of input parameters, data are observed at $m=20$ points over the time domain $[0,1]$ with $\xi=5\%$ relative noise. Five hundred posterior samples are used to generate ROM solutions, and the resulting distribution is characterized by the pointwise sample mean and $95\%$ interquantile range.
@@ 7 Concluding remarks
There are several avenues for future work related to GP-BayesOpInf. Our experiments used squared-exponential GP kernels and linear dimensionality reduction by POD, but the methodology is flexible and can be applied with any sufficiently differentiable GP kernel and with any type of compression (including nonlinear manifold approximators such as in [3, 22, 23, 34, 58]). It would be interesting to explore if embedding desirable properties in the kernel can improve model performance for problems with specific characteristics (for example, periodicity). In the presence of large amounts of noise, the number of reliable reduced modes that can be obtained through POD is limited. To reduce the projection error by including more reduced modes in GP-BayesOpInf, it may be advantageous to leverage filtering strategies as shown in [20] prior to applying dimensionality reduction. We also plan to explore more advanced GP techniques, such as sparse GPs [64, 65] to reduce the trajectory data needed for regression (i.e., selecting $m^{\prime}$ automatically), multifidelity GPs [2, 12, 40, 45, 49, 50] to fuse data from different sources, as well as a single vector-valued GP for all reduced modes to incorporate physically meaningful spatial correlations. Another promising direction is to leverage the quantified uncertainties to guide the active selection of initial conditions and/or input functions for parsimoniously generating informative training trajectories.